# Forecasting with Hyper-Trees

### v1

@@ 4.1 Local Time Series Models: Illustrative Example
All classical forecasting models are estimated using the Nixtla-implementation of Garza et al., (2022). We train the HyperTree-AR(12) and the HyperTree-ES, as well as all tree-based models using 500 iterations, a learning rate of $\eta=0.1$ and using the linear-tree option, without additional hyper-parameter tuning. [^17] We train the Hyper-Tree models with the linear-tree option. In contrast to the forecasts of the base LightGBM (Ke et al.,, 2017), however, forecasts of the Hyper-Tree models are not trend-flat even if the linear-trees are not used. This is a result of the fact that forecasts of the Hyper-Tree models are created via the inductive bias of the target models. Hence, while forecasts of the classical LightGBM (Ke et al.,, 2017) are piece-wise constant without the linear-tree option, it is the forecasted parameters of the Hyper-Tree that are piece-wise constant, so that when applied to the target models, forecasts are still reasonable. All models leveraging covariates are specified using time-derived features $\mathbf{x}_{{}_{t}}=\{month,quarter\}$. LightGBM and LightGBM-sktime further use a time-index as an additional feature. In the remainder of this manuscript, we evaluate point forecasts only. However, as we have shown in Section 2.2, our architecture allows us to also extend the Hyper-Tree models to a probabilistic setting. We leave the probabilistic extension to future versions of this manuscript. Returning to the evaluation of the models for the air-passenger dataset, Table 1 shows the comparison across the different models.
@@
Table 1 shows that the Hyper-Tree models are competitive, with the HyperTree-AR being the most accurate across all metrics, except for RMSE, where LightGBM shows the highest accuracy. In contrast, while being the classical analog due to its architectural similarity to the HyperTree-AR, the AR-X exhibits the lowest accuracy across all metrics. This underscores the significance of parameter learning and forecasting using a tree-based hyper-model, which distinguishes the HyperTree-AR from AR-X. The HyperTree-AR appears to more effectively capture the multiplicative behavior of the seasonal component of the series, likely due to its non-stationary AR-parameters that are forecasted as a function of time-related features $\mathbf{x}_{{}_{t}}=\{month,quarter\}$, and their interactions. Additionally, it is important to mention that while the AutoARIMA has been fine-tuned for the specific dataset (the selected model is an ARIMA(1,1,0)(0,1,0)[12]), the HyperTree-AR is a plain AR(12) model without additional complexity or fine tuning. Hence, it is likely that the accuracy of the HyperTree-AR can be further improved via, e.g., seasonal differencing and/or selecting a more appropriate lag structure. Even though the HyperTree-ES compares well across metrics, it is not as competitive as the HyperTree-AR. Yet, when compared to its classical analog, the HyperTree-ES outperforms the AutoETS model across all metrics. The HyperTree-ES also compares favorably well to the classical AutoARIMA and AutoARIMA-X models. This again highlights the benefits of using a Hyper-Tree architecture over their traditional counterparts. The forecasts of the HyperTree-AR and the second best LightGBM are visualized in Figure 8.
@@ 4.2 Local Time Series Models: Extended Evaluation
In the previous section, we applied and evaluated the Hyper-Tree architectures to a single, well behaved time series that exhibits strong seasonality and trend. As such, the repeating pattern of the series can easily be captured and modeled using time derived features, e.g., $\mathbf{x}_{{}_{t}}=\{month,quarter\}$, leading to excellent forecasting results. To further evaluate and compare the Hyper-Tree architecture across heterogeneous data characteristics, we extend the evaluation to the following datasets: [^18] We plan to extend the comparison with additional datasets in later versions of the manuscript. A natural choice is the M5 data (Makridakis et al., 2022a, ; Makridakis et al., 2022b, ).
@@
Table 2 describes the characteristics of each dataset.
Table 2: Dataset Description.
@@
We opted to use the Rossmann Store Sales over the M5 data since the former is on store rather than on article level, reducing the amount of intermittency in the data. This decision was influenced by the autoregressive nature of the models, which would produce primarily zero forecasts when applied to the most disaggregated level of the M5 data, due to fact that the parameters are multiplied with mostly zero-valued lags of the series. This would make it challenging to accurately evaluate the effectiveness across models. We use the same set of models as in the previous section, adjusting the lag-structure based on the frequency of the data. However, we exclude the HyperTree-ES from the evaluation. This decision stems mainly from the recursive nature of the exponential smoothing parts, which drastically increases runtime for the Hyper-Tree model, especially for long series as available in the Rossmann Store Sales dataset. [^19] We are currently working on a runtime efficient version of the HyperTree-ES, planning to include it in future versions of the manuscript. Table 3 shows the hyper-parameter settings used for the experiments.
@@
Dataset Iterations Learning Rate Linear-Tree Early-Stopping Loss Features
@@
  • All other LightGBM Hyper-Parameters are kept at their default values. The above settings are used for the Hyper-Trees, as well as for the LightGBM and LightGBM-sktime models. For the latter two, a time-index feature is used in addition to the ones reported. For the LightGBM-sktime model, we specify a cubic polynomial for de-trending and de-seasonalizing the series.
To present our results and to evaluate different models, we use boxplots in Figures 9, 10 and 11 instead of presenting the results in a table, since we believe that boxplots offer a more comprehensive picture of the error distribution than a single number. [^20] The y-axes in all boxplots are on the log10-scale and each plot is arranged by decreasing median values. Further results per dataset are available via https://github.com/StatMixedML/Hyper-Trees/tree/main/experiments.
@@
From the error metrics across datasets and models, we can infer that the Hyper-Tree model is competitive, both in terms of median accuracy, as well as in terms of variability as measured by the inter-quartile ranges. For the Australian Retail Turnover dataset, where most series are characterized by a pronounced trend and seasonality, the HyperTree-AR model is a close second behind the AutoETS. This confirms that Exponential Smoothing-type models (Holt,, 1957; Winters,, 1960) provide accurate and robust forecasts for seasonal data (Hyndman and Athanasopoulos,, 2021). [^21] The accuracy of the AutoETS underscores the need for a runtime-efficient implementation of the HyperTree-ES, enabling its inclusion in future comparisons. For the monthly Tourism dataset, the autoregressive structure of our Hyper-Tree model appears to be more adequate to capture the temporal pattern across series. Turning our attention to the more complex Rossmann Store Sales dataset, we see that our Hyper-Tree model performs exceptionally well across all error metrics, outperforming its classical counterparts. This is where the effectiveness of the modeling of the AR-parameters via the Hyper-Tree comes into play, since the series are less well-behaved and external factors influence their behavior in a non-linear manner. Again, it is interesting to note that while the AutoARIMA and AutoARIMA-X models are fine tuned for each series, the HyperTree-AR model tends to be more accurate, even though it is trained using a fixed set of hyper-parameters and lag-structure across all series. In contrast to the Australian Retail Turnover or the Tourism dataset, the AutoETS ranks last in the comparison for the Rossmann Store Sales dataset. We analyzed the AutoETS forecasts for several sample series and observed that they were generally flat, lacking to accurately represent the behavior of the series. In addition to the per-dataset boxplots presented in Figures 9, 10 and 11, Figure 12 presents boxplots across all datasets and series to facilitate model evaluation.
@@
Comparing the models across all datasets, Figure 12 shows that our Hyper-Tree architecture provides robust and accurate forecasts across a variety of heterogeneous time series. [^22] It should be noted that the Rossmann Store Sales data predominantly influences the boxplots due to its high number of series. Figure 13 reports runtimes for all models on the Rossmann Store Sales dataset, showing the HyperTree-AR model to have higher runtimes compared to most other models. [^23] Recall from Table 3 that all tree-based models are trained with 500 iterations so that the runtime is a function of boosting rounds. The higher runtime of the Hyper-Tree is a direct consequence of the one vs. all estimation strategy of tree-based models, where a separate tree is grown for each parameter of the target model. Since we use a lag-order of $p=21$ for the Rossmann Store Sales dataset, the HyperTree-AR model is around 21 times as slow as compared to the LightGBM model that only estimates a single parameter.
@@
In Figures 14, 15 and 16, we plot sample forecasts for the evaluated datasets, along with WAPE error measures in the legend. Note that we plot the HyperTree-AR model alongside the best benchmark model, so that there is a variation in the secondary model from one plot to another. [^24] In Figures 14, 15 and 16, the context length of the actuals is set to four times the forecast horizon. All forecast plots are available via https://github.com/StatMixedML/Hyper-Trees/tree/main/experiments.
@@
Figures 14, 15 and 16 show that the HyperTree-AR model effectively captures the underlying patterns of the series, demonstrating good accuracy across diverse datasets. Given that the HyperTree-AR model employs a uniform, non-optimized configuration across all series, which may compromise its forecast accuracy for certain series, optimizing the selection of hyper-parameters, features, and lag-structure can likely further enhance its accuracy.
@@ 5 Conclusion, Limitations, and Future Research
Looking forward, the horizon of future research in this domain is broad and inviting. With this version of the manuscript, we have only started to explore the effectiveness of tree-based hyper-models. As outlined in the text, we plan to extend the evaluation to a global setting, leveraging the cross-learning ability of trees to capture both inter- and intra-series patterns. Also, extending the evaluation to a broader set of datasets is on our agenda. The challenge of effectively capturing various temporal hierarchies and ensuring the robustness of the hyper-tree model across diverse time series and hyper-parameter settings remains a pivotal point of future investigation. Further, while this study has focused on training relatively simple target models, exploring the applicability of our approach to more sophisticated architectures poses an interesting extension. Leveraging the cross-learning abilities of tree-based models to allow for domain adaption or few-shot learning is another extension we are planning to incorporate into later versions of the manuscript. A probabilistic extension of our approach presents an additional compelling avenue for future research. As a limitation of our approach, it is important to note that most tree-based models employ a one vs. all estimation strategy, where a separate tree is grown for each parameter of the target model. This can substantially increase the computational burden, which in turn limits the scalability of our approach for target models with a high number of trainable parameters. Hence, it is crucial to develop more efficient tree structures that mitigate the issue of linear scaling with respect to the number of estimated parameters.

### v2

@@ 4.1 Local Time Series Models: Illustrative Example
All classical forecasting models are estimated using the Nixtla-implementation of Garza et al., (2022). We train the HyperTree-AR(12) and the HyperTree-ES, as well as all tree-based models using 500 iterations, a learning rate of $\eta=0.1$ and using the linear-tree option. Note that none of the hyper-parameters for the tree-based models in this manuscript are optimized. [^17] While we train the Hyper-Tree models with the linear-tree option, forecasts of the Hyper-Tree models are not trend-flat even if the linear-trees are not used, which does not hold true for the forecasts of the base LightGBM (Ke et al.,, 2017) without the linear-tree option. This is a result of the fact that forecasts of the Hyper-Tree models are created via the inductive bias of the target models. Hence, while forecasts of the classical LightGBM (Ke et al.,, 2017) are piece-wise constant without the linear-tree option, it is the forecasted parameters of the Hyper-Tree that are piece-wise constant, so that when applied to the target models, forecasts are still reasonable. All models leveraging covariates are specified using time-derived features $\mathbf{x}_{{}_{t}}=\{month,quarter\}$. LightGBM and LightGBM-sktime further use a time-index as an additional feature. In the remainder of this manuscript, we evaluate point forecasts only. [^18] However, as we have shown in Section 2.2, our architecture allows us to also extend the Hyper-Tree models to a probabilistic setting. We leave the probabilistic extension to future versions of this manuscript. Table 1 presents a comparison of the different models.
@@
Table 1 shows that the Hyper-Tree models are competitive, with the HyperTree-AR being the most accurate across all metrics, except for RMSE, where LightGBM shows the highest accuracy. In contrast, while being the classical analog due to its architectural similarity to the HyperTree-AR, the AR-X exhibits the lowest accuracy across all metrics. This underscores the significance of parameter learning and forecasting using a tree-based hyper-model, which distinguishes the HyperTree-AR from AR-X. The HyperTree-AR appears to more effectively capture the multiplicative behavior of the seasonal component of the series, likely due to its non-stationary AR-parameters that are forecasted as a function of time-related features $\mathbf{x}_{{}_{t}}=\{month,quarter\}$, and their interactions. Additionally, it is important to mention that while the AutoARIMA has been fine-tuned for the specific dataset (the selected model is an ARIMA(1,1,0)(0,1,0)[12]), the HyperTree-AR is a plain AR(12) model without additional complexity or fine-tuning. Hence, it is likely that the accuracy of the HyperTree-AR can be further improved via, e.g., seasonal differencing and/or selecting a more appropriate lag structure. Even though the HyperTree-ES compares well across metrics, it is not as competitive as the HyperTree-AR. Yet, when compared to its classical analog, the HyperTree-ES outperforms the AutoETS model across all metrics. The HyperTree-ES also compares favorably well to the classical AutoARIMA and AutoARIMA-X models. This again highlights the benefits of using a Hyper-Tree architecture over their traditional counterparts. The forecasts of the HyperTree-AR and the second best LightGBM are visualized in Figure 8.
@@ 4.2 Local Time Series Models: Extended Evaluation
In the previous section, we applied and evaluated the Hyper-Tree architectures to a single, well behaved time series that exhibits strong seasonality and trend. As such, the repeating pattern of the series can easily be captured and modeled using time derived features, e.g., $\mathbf{x}_{{}_{t}}=\{month,quarter\}$, leading to excellent forecasting results. To further evaluate and compare the Hyper-Tree architecture across heterogeneous data characteristics, we extend the evaluation to the following datasets: [^19] We plan to extend the comparison with additional datasets in later versions of the manuscript. A natural choice is the M5 data (Makridakis et al., 2022a, ; Makridakis et al., 2022b, ) on an aggregated level.
@@
Table 2 provides an overview of some key characteristics for each dataset, including the number of series, frequency, mean length of the series, number of lags used for AR($p$) models, as well as the forecast horizon.
Table 2: Dataset Descriptions.
@@
We opted to use the Rossmann Store Sales over the M5 data since the former is on store rather than on disaggregated article level, thereby reducing the amount of intermittency in the data. This decision was influenced by the autoregressive nature of the models, which would predominantly yield zero forecasts when applied to the most disaggregated level of the M5 data. This is because the parameters would be multiplied by lags of the series that are mostly zero-valued. As such, the amount of intermittency would make it challenging to accurately evaluate the accuracy across models. We use the same set of models as in the previous section, with the lag structure adjusted according to the data frequency. However, we exclude the HyperTree-ES from the evaluation. This decision stems mainly from the recursive nature of the exponential smoothing parts, which drastically increases runtime for the Hyper-Tree model, especially for long series as available in the Rossmann Store Sales dataset. [^20] We are currently working on a runtime efficient version of the HyperTree-ES, planning to include it in future versions of the manuscript. We welcome suggestions via https://github.com/StatMixedML/Hyper-Trees/discussions. Table 3 shows the hyper-parameter settings used for the experiments.
@@
Dataset Iterations Learning Rate Linear-Tree Early-Stopping Training-Loss Features $\mathbf{x}_{{}_{t}}$
@@
  • No hyper-parameter optimization is used. All LightGBM Hyper-Parameters, except the ones reported in this table, are kept at their default values. The above settings are used for the Hyper-Trees, as well as for the LightGBM, LightGBM-AR($p$) and LightGBM-sktime models. For LightGBM and LightGBM-sktime, a time-index feature is used in addition to the ones reported. For LightGBM-AR($p$), $\mathbf{x}_{{}_{t}}$ includes additional autoregressive lagged-target features. For LightGBM-sktime, we specify a cubic polynomial for de-trending and de-seasonalizing the series.
To present our results and to evaluate different models, we use boxplots in Figures 9, 10 and 11 instead of presenting the results in a table. We believe that boxplots offer a more comprehensive picture of the error distribution than a single number. [^21] The y-axes in all boxplots are on the log10-scale and models arranged by decreasing median values. For ease of visualization, we removed the top 1% highest forecast errors for each metric across all models but used all data to sort the models. More results are available via https://github.com/StatMixedML/Hyper-Trees/tree/main/experiments.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
From the error metrics across datasets and models, we can infer that the HyperTree-AR model is competitive, both in terms of accuracy, as well as in terms of variability as measured by the inter-quartile ranges. For the Australian Retail Turnover dataset, where most series are characterized by a pronounced trend and seasonality, the HyperTree-AR model performs well, closely trailing the AutoARIMA and AutoETS models. The evaluation results confirm that Exponential Smoothing-type models (Holt,, 1957; Winters,, 1960) provide accurate and robust forecasts for seasonal data (Hyndman and Athanasopoulos,, 2021). [^22] The accuracy of the AutoETS underscores the need for a runtime-efficient implementation of the HyperTree-ES, enabling its inclusion in future comparisons. For the monthly Tourism dataset, the autoregressive structure of our HyperTree-AR model appears to be more adequate to capture the temporal pattern across series. Turning our attention to the more complex Rossmann Store Sales dataset, we see that our HyperTree-AR model performs well across all error metrics, outperforming its classical counterparts. The benefit of modeling the AR-parameters via the Hyper-Tree is especially apparent in this context, as the series are less well-behaved and external factors potentially influence their behavior in a non-linear manner. It is interesting to note that while the AutoARIMA and AutoARIMA-X models are fine tuned for each series, the HyperTree-AR model tends to be more accurate, even though it is trained using a fixed set of hyper-parameters and lag-structure across all series. In contrast to the Australian Retail Turnover or the Tourism dataset, the AutoETS ranks last in the comparison for the Rossmann Store Sales dataset. We analyzed the AutoETS forecasts for several sample series and observed that they were generally flat, lacking to accurately represent the behavior of the series. In addition to the per-dataset boxplots presented in Figures 9, 10 and 11, Figure 12 presents boxplots across all datasets and series to facilitate model evaluation.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Comparing the models across all datasets, Figure 12 shows that our Hyper-Tree architecture provides robust and accurate forecasts across a variety of heterogeneous time series. [^23] It should be noted that the Rossmann Store Sales data predominantly influences the boxplots due to its high number of series. When contrasted with the LightGBM-AR model, which induces an autoregressive structure through its features and directly creates forecasts, our HyperTree-AR framework tends to show a higher accuracy across all datasets. Figure 13 reports runtimes for all models on the Rossmann Store Sales dataset, showing the HyperTree-AR model to have higher runtimes compared to most other models. [^24] Recall from Table 3 that all tree-based models are trained with 500 iterations so that the runtime is a function of boosting rounds. The higher runtime of the Hyper-Tree is a direct consequence of the one vs. all estimation strategy of tree-based models, where a separate tree is grown for each parameter of the target model. Since we use a lag-order of $p=21$ for the Rossmann Store Sales dataset, the HyperTree-AR model is around 21 times as slow as compared to the LightGBM model that only estimates a single parameter.
@@
In Figures 14, 15 and 16, we plot sample forecasts for the evaluated datasets, along with WAPE error measures in the legend. Note that we plot the HyperTree-AR model alongside the best benchmark model, so that there is a variation in the secondary model from one plot to another. [^25] In Figures 14, 15 and 16, the context length of the actuals is set to four times the forecast horizon.
@@
Figures 14, 15 and 16 show that the HyperTree-AR model effectively captures the underlying patterns of the series, demonstrating good accuracy across diverse datasets. Given that the HyperTree-AR model employs a uniform, non-optimized configuration across all series, which may compromise its forecast accuracy for certain series, optimizing the selection of hyper-parameters, features, and lag-structure can likely further enhance accuracy.
@@ 5 Conclusion, Limitations, and Future Research
Looking forward, the horizon of future research in this domain is broad and inviting. With this version of the manuscript, we have only started to explore the effectiveness of tree-based hyper-models. As outlined in the text, we plan to extend the evaluation to a global setting, leveraging the cross-learning ability of trees to capture both inter- and intra-series patterns. Also, extending the evaluation to a broader set of datasets is on our agenda. The challenge of effectively capturing various temporal hierarchies and ensuring the robustness of the hyper-tree model across diverse time series and hyper-parameter settings remains a pivotal point of future investigation. Further, while this study has focused on training relatively simple target models, exploring the applicability of our approach to more sophisticated architectures poses an interesting extension. Leveraging the cross-learning abilities of tree-based models to allow for domain adaption or few-shot learning is another extension we are planning to incorporate into later versions of the manuscript. A probabilistic extension of our approach presents an additional compelling avenue for future research. As a limitation of our approach, it is important to note that most tree-based models employ a one vs. all estimation strategy, where a separate tree is grown for each parameter of the target model. This can substantially increase the computational burden, which in turn limits the scalability of our approach for target models with a high number of trainable parameters. It is therefore crucial to develop more efficient tree structures that mitigate the issue of linear scaling with respect to the number of estimated parameters.